OpenStudy (anonymous):

The system, 6x - 4y = 5 and 9x - 6y = 5, is consistent. True False

OpenStudy (anonymous):

its false right?

OpenStudy (anonymous):

do you know what is consistent and inconsistent ?

OpenStudy (anonymous):

yes but u can describe it further :)

OpenStudy (anonymous):

but for that you have to describe how much you know first ? :)

OpenStudy (anonymous):

all i know is that its a system

OpenStudy (ash2326):

A system is consistent if it has atleast one solution. A system is inconsistent if it has no solutions.

OpenStudy (anonymous):

thx ash :)

OpenStudy (anonymous):

Just adding to that abstract definition it's really the concept of geometry of straight line in a plane.

OpenStudy (anonymous):

as far your question goes ..

OpenStudy (ash2326):

yeah, if you plot these two equations, you'll get two lines which don't intersect, that's why there is no solution if the lines intersect there is a solution

OpenStudy (anonymous):

what she meant is this two are parallel lines hence they would never intersect.

OpenStudy (anonymous):

i hate plottinglol

OpenStudy (anonymous):

but why ?

OpenStudy (ash2326):

not she, he :D

OpenStudy (anonymous):

Oh sorry :D

OpenStudy (ash2326):

no problem

OpenStudy (anonymous):

math was never my gift...:(

OpenStudy (mathteacher1729):

Change them to slope-intercept form: 6x - 4y = 5 add 6x to both sides -4y = -6x + 5 divide both sides by -4 y = (-6)/(-4)x - 5/4 simplify y = 3/2x - 5/4 Now the other equation: 9x - 6y = 5 subtract 9x from both sides -6y = -9x + 5 divide by -6 y = (-9)/(-6)x -5/6 simplify y = 3/2x -5/6 Now let's compare the SLOPES and the INTERCEPTS of y = 3/2x - 5/4 and y = 3/2x -5/6 They have the SAME slope (3/2) so they are PARALLEL they have DIFFERENT y-intercepts (-5/4) and (-5/6) respectively. So that means the lines are parallel and do not overlap. ================= The system is INCONSISTENT. ================= You can see this by going to http://www.wolframalpha.com and typing plot 6x - 4y = 5 and 9x - 6y = 5 Hope this helps!

OpenStudy (ash2326):

hey lisa math can be learnt by practice :)

OpenStudy (anonymous):

omg that really help!

OpenStudy (anonymous):

ash: this is my second year repeating algebra :(

OpenStudy (anonymous):

There are other tricky ways in linear algebra that could solve this task just by seeing the equation.

OpenStudy (anonymous):

for example just consider the ratio of the two coefficients they are equal but not equal to the ratio of the constants hence inconsistent.

OpenStudy (ash2326):

lisa: don't be upset, work hard and u'll pass with flying colors

OpenStudy (anonymous):

thx ash :)... the first semester is almost over and i dont even have half of my work done lol

OpenStudy (anonymous):

and again just try to find the determinant of the coefficient matrix you will see that this is a singular and hence the system is inconsistent.

OpenStudy (anonymous):

foolformath thx for the information <3

OpenStudy (anonymous):

how old are you lisa ?

OpenStudy (anonymous):

17

OpenStudy (anonymous):

you are welcome ..

OpenStudy (anonymous):

cool you have lots of time to learn mathematics :)

OpenStudy (ash2326):

lisa it's never too late to start right now

OpenStudy (anonymous):

thx guys :)... i just hope to graduate by next year...im home schooled..

OpenStudy (anonymous):

wow !! that's even better .. just remember that mathematics couldn't be learned unless you are in love with it .. it's a beautiful subject just need a little bit of you attention and care .. treat it like your boyfriend :P and it will love you back :D

OpenStudy (anonymous):

lol thats good advice :) haha

OpenStudy (anonymous):

hehe :D :)

OpenStudy (anonymous):

so r u a teacher?

OpenStudy (ash2326):

me?

OpenStudy (anonymous):

both of you :)

OpenStudy (ash2326):

no i'm just a graduate , working now, will go for post graduation in a year or so

OpenStudy (anonymous):

I am just another student who is a fool :P

OpenStudy (anonymous):

haha you guys are cool!....

OpenStudy (ash2326):

stay hungry stay foolish :D

OpenStudy (anonymous):

hehe :D but I love to eat btw :P

OpenStudy (anonymous):

Maths graduate ash ?

OpenStudy (ash2326):

nope , engineering graduate :D

OpenStudy (anonymous):

hahah who doesnt like to eat? but i recently became a vegetarian its hardlol

OpenStudy (anonymous):

duh not good :P.. what's up with you girls ?? My gf too become veggie :(

OpenStudy (ash2326):

that's good lisa, i'm also a vegetarian . never tasted meat or anything like that

OpenStudy (anonymous):

Veggie ?? how can you live without that? :P

OpenStudy (anonymous):

well i watched this show about meat and it was talkin about how meat is causeing cancer....and thats cool ash! :)

OpenStudy (anonymous):

well ur gf is a smart girl :)

OpenStudy (ash2326):

yeah ur gf is smart, u should learn from her

OpenStudy (anonymous):

Sorry, I can't give up meat .... my body and mind refuses to work without it .. :P

OpenStudy (anonymous):

I know she is smart.. how else I would be dating her ? :P hehe :D

OpenStudy (ash2326):

yeah :D

OpenStudy (anonymous):

lol she has a smart boy friend ;) but he needs to give up meat loljk

OpenStudy (anonymous):

i just broke up with my boy friend.. he was jerklol

OpenStudy (anonymous):

I am no near to be smart .. I am just a FOOL :P

OpenStudy (ash2326):

oh you'd fine a better guy soon :D

OpenStudy (anonymous):

and I am sorry hear that lisa . :(

OpenStudy (ash2326):

ffm ur being modest

OpenStudy (ash2326):

lisa and foolformath what's the time there

OpenStudy (anonymous):

:)...its ok....i finally realized i dont need him.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

I am not ash, you don't know how much I have to work to remember simple things :anybody who works like me would do much better .. I believe

OpenStudy (anonymous):

lisa you are 17 .. boys will come and boys will go .. do not worry :)

OpenStudy (anonymous):

ffm well whatever it take u still r smart :) and thx...i just hope i dont keep getting hurt... it sucks lol

OpenStudy (ash2326):

keep practicing you'd be able to remember more :)

OpenStudy (anonymous):

I know the feeling .. I was once dumped like that :P but you know then I met the most beautiful thing Mathematics .. now we are together forever :D

OpenStudy (anonymous):

easy to say than done :P

OpenStudy (ash2326):

yeah lisa, i also hope so, may u get a guy who cares for you the most

OpenStudy (anonymous):

lol aww u really do have a passion for math..thats awsome

OpenStudy (anonymous):

aww thx ash :)

OpenStudy (anonymous):

btw i didnt get dumped...lol

OpenStudy (ash2326):

fool for math that's very cool.

OpenStudy (anonymous):

oh yes only guys get dumped :P hearless girls ..... :D

OpenStudy (anonymous):

jokes :D

OpenStudy (anonymous):

lol i know :D

OpenStudy (ash2326):

i don't know that, never been dumped, with my girl since2 years

OpenStudy (anonymous):

wow ash thats cool! u two must really love eachother:)

OpenStudy (anonymous):

the longest iv been with a guy is 4 months

OpenStudy (anonymous):

wow excellent track record lisa :P

OpenStudy (ash2326):

that happens in this age, u enjoy and live it to the fullest, won't get this time again

OpenStudy (anonymous):

lol ;)

OpenStudy (anonymous):

yea but i wish i was older already

OpenStudy (ash2326):

you know i'd also wished this, but now i feel that time was much better

OpenStudy (anonymous):

lol yea... btw i miss public school...but i got in trouble alote.....lets just say the wrong crowd

OpenStudy (anonymous):

public school ? it's good ? ??

OpenStudy (anonymous):

well i liked it cause i loved being wit my friends everyday.... now i barely see them

OpenStudy (ash2326):

oh lisa that's sad:(

OpenStudy (anonymous):

that's sad ..

OpenStudy (ash2326):

hey foolfor math what's ur name

OpenStudy (anonymous):

I have never studied in public school myself so I really don't know but they say it's not good..

OpenStudy (anonymous):

I am Max :)

OpenStudy (ash2326):

cool name

OpenStudy (anonymous):

My full name is Maximus Decimus Meridius! guess who I am :P

OpenStudy (anonymous):

haahah :D..